# Mix-of-Granularity: Optimize the Chunking Granularity for Retrieval-Augmented Generation

### v1

@@ A.4 PubMedQA*
A.5 BioASQ-Y/N
BioASQ [47] is an annual biomedical QA competition. For the MIRAGE benchmark, we selected only the Machine Reading Comprehensive Track (Task B), focusing on 618 questions from recent years (2019-2023).
Appendix B Exact versions of LLMs
@@
Appendix C Experiment on soft labels
In this section, we evaluate the performance of building soft labels using different methods (TF-IDF [38], RoBERTa [32], and hitrate score). For this experiment, we fix one retriever (BM25) and the backbone LLM (Qwen1.5), then build the soft labels using these three different methods. Different routers are trained with these different soft labels and tested on various datasets of the Medical Question-Answering task. The results are grouped in Table 5.
The results indicate that there is no universal best method for building soft labels. For instance, “hitrate score” is the most effective when training on MedMCQA, while RoBERTa shows advantages on PubMedQA. Overall, the soft labels built with RoBERTa demonstrate good performance across the board. Therefore, for the remainder of the experiments, RoBERTa is adopted as the default method for building soft labels.
@@ Appendix C Experiment on soft labels
Appendix D Impact of the Number of Candidate Snippets
@@
Appendix E Choice of Retriever
In the previous experiment, BM25 [39] was used as the retriever because it is a lightweight and popular choice in practice. In this section, we replace BM25 with other popular retrievers to evaluate their performance. Referring to the setup in MedRAG [51], we select a general-domain semantic retriever called Contriever [18], a scientific-domain retriever called SPECTER [7], and a biomedical-domain retriever called MedCPT [23]. Additionally, the Reciprocal Rank Fusion (RRF) method [9] is utilized to combine results from different retrievers, including RRF-2 (fusion of BM25 and MedCPT) and RRF-4 (fusion of all four retrievers). From Table 6, we can observe the following: Each retriever has a well-performing QA set, which results in minimal overall differences among these retrievers. Considering these conclusions, we decide to continue using BM25 for all other experiments.
Table 6: Performance of different retrievers
@@ Appendix E Choice of Retriever
Appendix F Performance of MoG on Medical QA Task (Other results)
@@
Appendix G Analysis on Samples Improved or Degraded
@@
Appendix H Performance of MoGG on Medical QA Task (Other results)
@@
Table 8: Accuracy of Medical Question-Answering task with MoGG (trained with StatPearls), best results marked in bold.

### v2

@@ B.4 PubMedQA*
B.5 BioASQ-Y/N
BioASQ Tsatsaronis et al. (2015) is an annual biomedical QA competition. For the MIRAGE benchmark, we selected only the Machine Reading Comprehensive Track (Task B), focusing on 618 questions from recent years (2019-2023).
The important statistics of the copora and the QA datasets are presente in Table 3 below.
Corpus #Doc #Snip. $\bar{L}$. Domain
PubMed 23.9M 23.9M 196 Bio.-Med.
StatPearls 9.3k 301.2k 119 Clinics
Textbooks 18 125.8k 182 Medicine
Wikipedia 6.5M 29.9M 162 General
MedCorp 30.4M 54.2M 221 Mixed
Datasets Size #Opt. $\bar{L}$ Source
MMLU-Med 1,089 4 63 Exam.
MedQA-US 1,273 4 177 Exam.
MedMCQA 4,183 4 26 Exam.
PubMedQA* 500 3 24 Literature
BioASQ-Y/N 618 2 17 Literature
Table 3: Important statistics of corpora and QA datasets
Appendix C Exact Versions of LLMs
@@
Table 4: Versions of the backbone LLMs
Appendix D Qualitative Results of the Router
As illustrated in Figure 4, the router effectively assigns different weights to various granularity levels. From the figure, we can infer that a potential global peak in weights may exist at a granularity level smaller than level 1 or larger than level 5.
However, we did not test these smaller or larger granularity levels due to the following reasons: In our experiments, we focus exclusively on LLMs with a parameter size of around 7 billion. For these models, the chunking size at granularity level 5 approaches their maximum context window. Conversely, the chunking size at granularity level 1 consists of only a few dozen characters, which is already quite small.
Appendix E Experiment on Soft Labels
In this section, we evaluate the performance of building soft labels using different methods (TF-IDF Ramos (2003), RoBERTa Liu et al. (2019), and hitrate score). For this experiment, we fix one retriever (BM25) and the backbone LLM (Qwen1.5), then build the soft labels using these three different methods. Different routers are trained with these different soft labels and tested on various MQA datasets. The results are grouped in Table 5.
The results indicate that there is no universal best method for building soft labels. For instance, “hitrate score” is the most effective when training on MedMCQA, while RoBERTa shows advantages on PubMedQA. Overall, the soft labels built with RoBERTa demonstrate good performance across the board. Therefore, for the remainder of the experiments, RoBERTa is adopted as the default method to build soft labels.
@@ Appendix E Experiment on Soft Labels
Appendix F Impact of the Number of Candidate Snippets
@@
Appendix G Choice of Retriever
In the previous experiment, BM25 Robertson and Zaragoza (2009) was used as the retriever because it is a lightweight and popular choice in practice. In this section, we replace BM25 with other popular retrievers to evaluate their performance. Referring to the setup in MedRAG Xiong et al. (2024), we select a general-domain semantic retriever called Contriever Izacard et al. (2022a), a scientific-domain retriever called SPECTER Cohan et al. (2020), and a biomedical-domain retriever called MedCPT Jin et al. (2023). Additionally, the Reciprocal Rank Fusion (RRF) method Cormack et al. (2009) is utilized to combine results from different retrievers, including RRF-2 (fusion of BM25 and MedCPT) and RRF-4 (fusion of all four retrievers). From Table 6, we can observe the following: Each retriever has a well-performing QA set, which results in minimal overall differences among these retrievers. Considering these conclusions, we decide to continue using BM25 for all other experiments.
@@ Appendix G Choice of Retriever
Table 6: Performance of different retrievers
Appendix H Performance of MoG on Medical QA Task (Other results)
@@
Appendix I Analysis on Samples Improved or Degraded
@@
Appendix J Performance of MoGG on Medical QA Task (Other results)
@@
Table 8: Accuracy of Medical Question-Answering task with MoGG (trained with StatPearls), best results marked in bold
Appendix K MoGG’s Abaltion test
In this section, we rerun the MoGG without the router on Llama 3 to highlight its effectiveness. The results of MoGG’s ablation test is in Table 9 (highlighted in bold). The results are similar to MedRAG, because, without the router, MoGG is essentially the same as MedRAG, the added flexibility of the graph will only show when used with a router.
Method MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
CoT 0.7079 0.6431 0.5663 0.5500 0.7258 0.6386
MedRAG 0.6485 0.5961 0.4146 0.3800 0.5242 0.5127
MoG 0.7228 0.6196 0.5484 0.5100 0.7097 0.6221
MoGG w/o router 0.6683 0.5451 0.4182 0.3800 0.5484 0.5120
MoGG 0.7070 0.5961 0.5460 0.5200 0.7661 0.6262
Table 9: Ablation test with Llama3 as backbone LLM
Appendix L MoG Trained With Only Textbooks Corpus
In this Appendix, we present the results of training MoG only with Textbooks corpus. As Textbooks are only a subset of MedCorp, the performance is not as good as the ones shown in Table 1. Compared with Table 10 and Table 2, we can see the advantage of MoGG more clearly.
LLM Method MIRAGE Benchmark Dataset (Acc.)
MMLU MedQA MedMCQA PubMedQA BioASQ Avg.
GLM3 MedRAG 0.4802 0.3569 0.3811 0.3600 0.5565 0.4269
MoG 0.5545 0.2941 0.3548 0.4700 0.5726 0.4492
GPT-3.5 MedRAG 0.7277 0.6745 0.4468 0.2600 0.5161 0.5250
MoG 0.7525 0.6667 0.5603 0.5200 0.7016 0.6122
InternLM MedRAG 0.6188 0.5569 0.3118 0.1400 0.4839 0.4223
MoG 0.7277 0.5725 0.5281 0.3400 0.7339 0.5804
Llama3 MedRAG 0.6485 0.5961 0.4146 0.3800 0.5242 0.5127
MoG 0.7228 0.6196 0.5484 0.5100 0.7097 0.6221
Qwen1.5 MedRAG 0.5941 0.4000 0.3835 0.3300 0.4919 0.4399
MoG 0.5792 0.3843 0.4110 0.3300 0.6129 0.4635
Table 10: Accuracy of Medical Question-Answering task with MoG (trained with only Textbooks)
Appendix M Detailed Analysis of Time Consumption
In this appendix, we present the detailed experiment results of execution time. We measure the wall-clock execution time with different number of granularities using Llama3 as the backbone LLM. The results organized in Table 11 show that:
(1) the introduction of the router module will effectively increase the inference time by about 60%;
(2) increasing the number of granularity levels will only increase a marginal increase in execution time. The reason is that the bottleneck of execution time is the API calling time of backbone LLMs.
# (Granularity) MMLU MedQA MedMCQA PubMedQA BioASQ Global Avg.
1 w/o router 6.61 9.02 10.01 7.40 7.53 8.12
1 w/ router 10.15 10.76 15.81 15.99 13.09 13.16
2 13.25 12.60 20.53 17.30 14.98 15.73
3 16.30 16.01 19.03 16.73 17.12 17.04
4 13.43 20.65 22.72 18.93 14.05 17.96
5 13.20 20.43 23.00 19.36 13.66 17.93
Table 11: Performance Metrics by Granularity